# Elevating All Zero-Shot Sketch-Based Image Retrieval Through Multimodal Prompt Learning

### v1

@@ Abstract
We address the challenges inherent in sketch-based image retrieval (SBIR) across various settings, including zero-shot SBIR, generalized zero-shot SBIR, and fine-grained zero-shot SBIR, by leveraging the vision-language foundation model, CLIP. While recent endeavors have employed CLIP to enhance SBIR, these approaches predominantly follow uni-modal prompt processing and overlook to fully exploit CLIP’s integrated visual and textual capabilities. To bridge this gap, we introduce SpLIP, a novel multi-modal prompt learning scheme designed to operate effectively with frozen CLIP backbones. We diverge from existing multi-modal prompting methods that either treat visual and textual prompts independently or integrate them in a limited fashion, leading to suboptimal generalization. SpLIP implements a bi-directional prompt-sharing strategy that enables mutual knowledge exchange between CLIP’s visual and textual encoders, fostering a more cohesive and synergistic prompt processing mechanism that significantly reduces the semantic gap between the sketch and photo embeddings. In addition to pioneering multi-modal prompt learning, we propose two innovative strategies for further refining the embedding space. The first is an adaptive margin generation for the sketch-photo triplet loss, regulated by CLIP’s class textual embeddings. The second introduces a novel task, termed conditional cross-modal jigsaw, aimed at enhancing fine-grained sketch-photo alignment, by focusing on implicitly modelling the viable patch arrangement of sketches using knowledge of unshuffled photos. Our comprehensive experimental evaluations across multiple benchmarks demonstrate the superior performance of SpLIP in all three SBIR scenarios. Code is available at https://github.com/mainaksingha01/SpLIP.
@@
Figure 1: (a, b) The difference between the existing multi-modal prompt learning (MPL) vs ours. As opposed to the literature [20, 21, 37], we propose to enhance the generalizability of both the textual and visual prompt embeddings with mutual knowledge sharing in a principled layer-wise fashion. (c, d) Proposed conditional cross-modal jigsaw vs the literature [39]. As against a triplet loss connecting sketch-photo with same permutation while contrasting against the photo with a different permutation, we propose better learning of patch arrangements by associating a permuted sketch with its intact photo counterpart positively through a novel objective.
@@ 1 Introduction
Hand-drawn sketches, with their simple yet evocative lines, adeptly convey abstract ideas. The advent of touchscreen mobile devices has propelled sketch-based image retrieval (SBIR) [25, 62] into the limelight, offering myriad practical uses. SBIR retrieves photos from a vast repository based on the same category as a query sketch. However, despite representing the same class, sketches and photos often differ noticeably due to their distinct domains. Traditional methods [14, 43] address domain heterogeneity, assuming full visibility of test classes during training, showing promise in effective retrieval. Yet, a more realistic challenge emerges when test set categories remain unseen during training, termed zero-shot SBIR (ZS-SBIR) [11, 5, 28]. Contrarily, generalized ZS-SBIR (GZS-SBIR) [11] encompasses retrieval photos of known and novel classes for novel-class sketch queries during inference, heightening complexity. Additionally, instance-level fine-grained ZS-SBIR (FG-ZS-SBIR) [58, 31] focuses on precise shape matching, intensifying challenges compared to category-level SBIR.
@@
Leading SBIR frameworks leveraging ConvNets and ViTs [5, 11, 52, 55] face intrinsic semantic constraints by their architecture. On the contrary, the advent of multi-modal foundational models, notably CLIP [34] and Align [19], has markedly enhanced visual comprehension by integrating visual and textual information. Though these VLMs have excelled in cross-domain learning, their SBIR integration remains limited. Initiatives like [39] have aimed to tailor CLIP to ZS-SBIR and FG-ZS-SBIR, focusing on visual prompts and patch shuffling to align sketches and photos at both micro and macro levels. Other efforts [29, 59] seek to finetune CLIP’s embeddings for SBIR, indicating an increasing desire to leverage these multi-modal platforms beyond original purposes.
Highlighting research gaps: Despite the success, these strategies often rely on simplistic, one-dimensional prompting, failing to fully harness CLIP’s dual-pathway synergy. This overlooks the rich, complementary insights of CLIP’s visual-textual fusion, leading to suboptimal ZS-SBIR performance. Addressing this, there’s an urgent call for novel approaches that dynamically utilize this combined knowledge to surpass the semantic limitations of existing models, thus expanding SBIR potentials.
@@
Proposed solution: Our proposed model, SpLIP, tackles these challenges by implementing a novel deep multimodal prompting approach (Fig. 1), facilitating efficient knowledge exchange between text and image branches of frozen CLIP. In CLIP’s text encoder, we enhance static textual tokens by incorporating additional latent visual tokens at each layer, departing from previous random initialization methods [20, 37]. Likewise, we enrich the image embeddings within the CLIP vision backbone with information from these image-conditioned textual token embeddings, summarized over all the semantic categories, layer by layer. This bidirectional information sharing mitigates the semantic gap in the obtained embeddings, thus contributing positively towards zero-shot inference.
Going forward, we introduce a novel conditional cross-modal jigsaw task aimed at strengthening the linkage between photo and sketch pairs for all the considered SBIR variants, predominantly for FG-ZS-SBIR. This method requires solving a jigsaw puzzle [30] using an anchor sketch, assisted by a positive and a negative image from the alternate modality while ensuring that the positive image significantly aids in this process, thereby enhancing model generalization through the learning of patch arrangement for reconstructing a complete image. This approach deviates from previous works [39, 33] that either deal with mixed-modal images for a naive jigsaw solver or employ triplet objectives with uniform or varying patch permutations for identifying positive and negative pairs, thus failing to bridge the gap between local and global contexts effectively (Fig. 1). Finally, we integrate the gold-standard cross-modal triplet loss for SBIR and a sketch/photo-text-based classification loss of CLIP, introducing an adaptive margin scheme for the triplet objective derived from the semantic class-prompt embeddings. Our salient contributions are, therefore,
@@
- Moreover, we enhance the conventional cross-modal triplet loss objective in ZS-SBIR by incorporating an adaptive margin scheme, leveraging CLIP’s textual prompt embeddings. Additionally, we introduce a novel conditional cross-modal jigsaw task aimed at refining the association between sketch and photo pairs.
- We conduct extensive experiments on three benchmark datasets: Sketchy-extended [27, 44], TUBerlin-extended [13], and QuickDraw-extended [7], covering (G)ZS-SBIR and FG-ZS-SBIR settings. SpLIP consistently outperforms existing competitors, achieving significant improvements in all the metrics.
@@ 2.1 Sketch-based image retrieval (SBIR)
(Generalized) Zero-shot SBIR: SBIR tasks involve retrieving photos corresponding to specific categories from a diverse collection of multi-category images based on a given sketch query, demanding a thorough understanding of the joint sketch-photo manifold. The literature is rich in fully-supervised SBIR endeavors utilizing deep and hand-crafted descriptors and involving different learning mechanisms including both generative and discriminative approaches [44, 18, 66, 50, 32, 43, 24]. Recently, [4] introduced a data-free training strategy for SBIR, relaxing the need for curating sketch-photo pairs.
Researchers have addressed challenges in recognizing unseen test-time classes through ZS-SBIR, facilitating category-level generalization. Extending the traditional Convnet-based frameworks [7, 11, 57, 28], graph convolutional networks, ViTs, and their combinations have been introduced to learn an unbiased shared feature space [26, 16, 63]. Recent advancements [39, 59, 29, 8] leverage CLIP’s zero-shot inference by integrating text with sketches and photos, outperforming counterparts. A combined loss function of supervised cross-entropy and metric objectives fosters a discriminative embedding space. On the other hand, GZS-SBIR permits the presence of training and test time photos during inference, causing the model to show high bias towards the training data, which has been tackled in the literature from different perspectives [11, 5, 67, 36, 26, 15, 29]. We take a different route to tackle (G/FG)ZS-SBIR through the notion of multi-modal prompting in CLIP, thus modeling the visual-semantic synergy effectively.
@@ 2.2 Vision-language models and multi-modal prompt learning
Recent research [65, 64, 2, 3, 48, 47, 49] highlights prompt learning as a viable alternative to VLM fine-tuning on downstream tasks. To this end, multi-modal deep prompts, unlike uni-modal approaches for CLIP, synergize its visual and textual components. [20] proposed to learn both visual and textual prompts, and showed that initializing the visual prompts from the textual counterparts enhances the performance. In addition, [21] and [37] introduced regularization and feature consistency to prevent overfitting and ensure textual variety. However, existing methods mainly employ unidirectional token sharing, restricting the overall generalizability and semantic depth of the learned embeddings. Contrarily, we propose a bilateral approach to disseminate relevant cross-modal insights across CLIP’s branches, establishing a more effective multi-modal prompting paradigm.
@@ 3 Methodology
In the following, we delve into the initialization of text inputs and provide a detailed explanation of the image-text embeddings of CLIP in Section 3.1. Moving forward, our image-driven textual prompting approach is elaborated upon in Section 3.2, while Section 3.3 addresses the proposed visual prompting methodology. For fine-grained sketch-photo feature association, we discuss the proposed conditional cross-modality jigsaw task and the related details in Section 3.4. Finally, in Section 3.5, we mention the considered loss objectives. A list of important variables used is summarized in the Supplementary.
@@ 3.1 Initialization of visual-textual embeddings
The pre-trained CLIP model operates on two modalities: text and image. It consists of a transformer-based [53] text encoder ($\mathcal{F}_{t}$) and ViT-based [9] image encoder ($\mathcal{F}_{v}$). Both of them consist of $\mathcal{L}$ transformer encoder layers. $\mathcal{F}_{t}$ generates feature representations for text descriptions to capture semantic information. Initially, it tokenizes the text inputs, consisting of $\mathcal{J}$ words, and projects them into word embeddings $\mathbf{W}_{0}=[w^{1}_{0},w^{2}_{0},\cdots,w^{\mathcal{M}}_{0}]\in\mathbb{R}^{\mathcal{M}\times d_{t}}$, where [.,.][.,.][ . , . ] denotes stacking and concatenation, $\mathcal{M}$ refers the number of embedding tokens ($77$ per class-prompt) and $d_{t}$ is the dimension of text tokens. In our approach, we use the input texts as ‘‘sketch/photo of a [CLS]’’ for sketches and photos, respectively, for (G)ZS-SBIR, meaning $\mathcal{J}=5$. Here, the [CLS] token represents class embeddings, completing the prompt embedding $\mathbf{W}_{0}$. For the FG-ZS-SBIR task, we use common text input, ‘‘visual representation of [CLS]’’, for both sketches and photos. We obtain the $l$-th layer embedding of $\mathcal{F}_{t}$, denoted as $\mathcal{F}_{t}^{l}$, as follows,
@@
$[\mathbf{c}_{l},\mathbf{E}_{l}]=\mathcal{F}_{v}^{l}([\mathbf{c}_{l-1},\mathbf{E}_{l-1}])\in\mathbb{R}^{1+\mathcal{N}\times d_{v}}\qquad l=1,2,\cdots,\mathcal{L}$ (2)
@@ 3.2 Proposed vision-guided deep textual prompting
[_$[\_,\mathbf{W}_{l}]=\mathcal{F}_{t}^{l}([\mathbf{T}_{l-1};\mathbf{W}_{l-1}])\in\mathbb{R}^{\mathcal{M}\times d_{t}}\qquad l=1,2,\cdots,\mathcal{L}$ (3)
@@ 3.3 Proposed text-guided deep visual prompting
In addition, we enforce token sharing from each layer of $\mathcal{F}_{t}$ to the respective layer input of $\mathcal{F}_{v}$. Note that, from $l\geq 1$, $\mathbf{W}_{l}$ implicitly includes visual knowledge as per our proposed textual prompting. Secondly, as opposed to the visual prompting approach proposed in [20], which only shares the learnable textual tokens with the visual branch, we propose to consider all the tokens from $\mathbf{W}_{l}$ over all the classes present in $\mathcal{C}^{s}$ to be included into the input space of $\mathcal{F}_{v}^{l}$. This class-agnostic knowledge-sharing approach helps better diminish the semantic gap in the learned embeddings.
@@
$[\mathbf{c}_{l},\_;\mathbf{E}_{l}]=\mathcal{F}_{v}^{l}([\mathbf{c}_{l-1},\mathbf{V}_{l-1};\mathbf{E}_{l-1}])\in\mathbb{R}^{1+\mathcal{N}\times d_{v}}\qquad l=1,2,\cdots,\mathcal{L}$ (5)
@@ 3.4 Proposed conditional cross-modal jigsaw solver for fine-grained sketch-photo feature association
To address these challenges, our proposed method incorporates positive and negative counterparts ($p_{a}^{+},p_{a}^{-}$) from set $\mathcal{P}^{s}$ when resolving the jigsaw puzzle for a sample $s_{a}\in\mathcal{S}^{s}$ given its permuted version $s_{a}^{\prime}$ which is obtained through a permutation function $\delta$, given the random permutation $y^{\text{perm}}\in\mathcal{Y}^{\text{perm}}$: $s^{\prime}_{a}=\delta(s_{a},y^{\text{perm}})$. Here, the information embedded in $(s^{\prime}_{a},p_{a}^{+})$ and $(s^{\prime}_{a},p_{a}^{-})$ is then processed through a transformer-based jigsaw-solver network, $\mathcal{F}_{js}$, operating atop $\mathcal{F}_{v}$, to leverage $p_{a}^{+}$ to better inform the prediction of the permutation arrangements of $s^{\prime}_{a}$ compared to $p_{a}^{-}$, achieved through a hinge objective. Collectively, $\mathcal{F}_{js}$ directly resolve the jigsaw puzzle for $s^{\prime}_{a}$ when coupled with $s_{a}$. These combined losses are referred to as $\mathcal{L}_{cjs}$. For simplicity, we define the fused inputs from pairs of images to $\mathcal{F}_{js}$ as follows: $r=[\mathcal{F}_{v}(s_{a}),\mathcal{F}_{v}(s^{\prime}_{a})]$, $r^{+}=[\mathcal{F}_{v}(p_{a}^{+}),\mathcal{F}_{v}(s^{\prime}_{a})]$, and $r^{-}=[\mathcal{F}_{v}(p_{a}^{-}),\mathcal{F}_{v}(s^{\prime}_{a})]$, respectively.

### v2

@@ Abstract
We address the challenges inherent in sketch-based image retrieval (SBIR) across various settings, including zero-shot SBIR, generalized zero-shot SBIR, and fine-grained zero-shot SBIR, by leveraging the vision-language foundation model CLIP. While recent endeavors have employed CLIP to enhance SBIR, these approaches predominantly follow uni-modal prompt processing and overlook to exploit CLIP’s integrated visual and textual capabilities fully. To bridge this gap, we introduce SpLIP, a novel multi-modal prompt learning scheme designed to operate effectively with frozen CLIP backbones. We diverge from existing multi-modal prompting methods that treat visual and textual prompts independently or integrate them in a limited fashion, leading to suboptimal generalization. SpLIP implements a bi-directional prompt-sharing strategy that enables mutual knowledge exchange between CLIP’s visual and textual encoders, fostering a more cohesive and synergistic prompt processing mechanism that significantly reduces the semantic gap between the sketch and photo embeddings. In addition to pioneering multi-modal prompt learning, we propose two innovative strategies for further refining the embedding space. The first is an adaptive margin generation for the sketch-photo triplet loss, regulated by CLIP’s class textual embeddings. The second introduces a novel task, termed conditional cross-modal jigsaw, aimed at enhancing fine-grained sketch-photo alignment by implicitly modeling sketches’ viable patch arrangement using knowledge of unshuffled photos. Our comprehensive experimental evaluations across multiple benchmarks demonstrate the superior performance of SpLIP in all three SBIR scenarios. Project page: https://mainaksingha01.github.io/SpLIP/.
@@
Figure 1: (a, b) The difference between the existing multi-modal prompt learning (MPL) vs ours. As opposed to the literature [20, 21, 37], we propose to enhance the generalizability of both the textual and visual prompt embeddings with mutual knowledge sharing in a principled layer-wise fashion. (c, d) Proposed conditional cross-modal jigsaw vs the literature [39]. As against a triplet loss connecting sketch-photo with the same permutation while contrasting against the photo with a different permutation, we propose better learning of patch arrangements by positively associating a permuted sketch with its intact photo counterpart through a novel objective.
@@ 1 Introduction
Hand-drawn sketches adeptly convey abstract ideas with their simple yet evocative lines. The advent of touchscreen mobile devices has propelled sketch-based image retrieval (SBIR) [25, 61] into the limelight, offering myriad practical uses. SBIR retrieves photos from a vast repository based on the same category as a query sketch. However, despite representing the same class, sketches and photos often differ noticeably due to their distinct domains. Traditional methods [14, 43] address domain heterogeneity, assuming full visibility of test classes during training, showing promise in effective retrieval. Yet, a more realistic challenge emerges when test set categories remain unseen during training, termed zero-shot SBIR (ZS-SBIR) [11, 5, 28]. Contrarily, generalized ZS-SBIR (GZS-SBIR) [11] encompasses retrieval photos of known and novel classes for novel-class sketch queries during inference, heightening complexity. Additionally, instance-level fine-grained ZS-SBIR (FG-ZS-SBIR) [58, 31] focuses on precise shape matching, intensifying challenges compared to category-level SBIR.
@@
Leading SBIR frameworks leveraging ConvNets and ViTs [5, 11, 52, 55] face intrinsic semantic constraints by their architecture. On the contrary, the advent of multi-modal foundational models, notably CLIP [34] and Align [19], has markedly enhanced visual comprehension by integrating visual and textual information. These VLMs have excelled in cross-domain learning, but their SBIR integration remains limited. Initiatives like [39] have aimed to tailor CLIP to ZS-SBIR and FG-ZS-SBIR, focusing on visual prompts and patch shuffling to align sketches and photos at both micro and macro levels. Other efforts [29, 59] seek to finetune CLIP’s embeddings for SBIR, indicating an increasing desire to leverage these multi-modal platforms beyond original purposes.
Highlighting research gaps: Despite the success, these strategies often rely on simplistic, one-dimensional prompting, failing to harness CLIP’s dual-pathway synergy fully. This overlooks CLIP’s visual-textual fusion’s rich, complementary insights, leading to suboptimal ZS-SBIR performance. Addressing this, there’s an urgent call for novel approaches that dynamically utilize this combined knowledge to surpass the semantic limitations of existing models, thus expanding SBIR potentials.
@@
Proposed solution: Our proposed model, SpLIP, tackles these challenges by implementing a novel deep multimodal prompting approach (Fig. 1), facilitating efficient knowledge exchange between text and image branches of frozen CLIP. In CLIP’s text encoder, we enhance static textual tokens by incorporating additional latent visual tokens at each layer, departing from previous random initialization methods [20, 37]. Likewise, we enrich the image embeddings within the CLIP vision backbone with information from these image-conditioned textual token embeddings, summarized layer by layer over all the semantic categories. This bidirectional information sharing mitigates the semantic gap in the obtained embeddings, thus contributing positively towards zero-shot inference.
As we advance, we introduce a novel conditional cross-modal jigsaw task to strengthen the linkage between photo and sketch pairs for all the considered SBIR variants, predominantly for FG-ZS-SBIR. This method requires solving a jigsaw puzzle [30] using an anchor sketch, assisted by a positive and a negative image from the alternate modality while ensuring that the positive image significantly aids in this process, thereby enhancing model generalization through the learning of patch arrangement for reconstructing a complete image. This approach deviates from previous works [39, 33] that either deal with mixed-modal images for a naive jigsaw solver or employ triplet objectives with uniform or varying patch permutations for identifying positive and negative pairs, thus failing to bridge the gap between local and global contexts effectively (Fig. 1). Finally, we integrate the gold-standard cross-modal triplet loss for SBIR and a sketch/photo-text-based classification loss of CLIP, introducing an adaptive margin scheme for the triplet objective derived from the semantic class-prompt embeddings. Our salient contributions are, therefore,
@@
- We enhance the conventional cross-modal triplet loss objective in ZS-SBIR by incorporating an adaptive margin scheme, leveraging CLIP’s textual prompt embeddings. Additionally, we introduce a novel conditional cross-modal jigsaw task aimed at refining the association between sketch and photo pairs.
- We conduct extensive experiments on three benchmark datasets: Sketchy-Ext [44, 11, 57], TU-Berlin-Ext [13, 27], and QuickDraw-Ext [17, 7], covering (G)ZS-SBIR and FG-ZS-SBIR settings. SpLIP consistently outperforms existing competitors, achieving significant improvements in all the metrics.
@@ 2.1 Sketch-based image retrieval (SBIR)
(Generalized) Zero-shot SBIR: SBIR tasks involve retrieving photos corresponding to specific categories from a diverse collection of multi-category images based on a given sketch query, demanding a thorough understanding of the joint sketch-photo manifold. The literature is rich in fully-supervised SBIR endeavors utilizing deep and hand-crafted descriptors and involving different learning mechanisms, including both generative and discriminative approaches [44, 18, 65, 50, 32, 43, 24]. Recently, [4] introduced a data-free training strategy for SBIR, relaxing the need for curating sketch-photo pairs.
Researchers have addressed challenges in recognizing unseen test-time classes through ZS-SBIR, facilitating category-level generalization. Extending the traditional ConvNet-based frameworks [7, 11, 57, 28], graph convolutional networks, ViTs, and their combinations have been introduced to learn an unbiased shared feature space [26, 16, 62]. Recent advancements [39, 59, 29, 8] leverage CLIP’s zero-shot inference by integrating text with sketches and photos, outperforming counterparts. A combined loss function of supervised cross-entropy and metric objectives fosters a discriminative embedding space. On the other hand, GZS-SBIR permits the presence of training and test time photos during inference, causing the model to show high bias towards the training data, which has been tackled in the literature from different perspectives [11, 5, 66, 36, 26, 15, 29]. We take a different route to tackle (G/FG)ZS-SBIR through multi-modal prompting in CLIP, thus modeling the visual-semantic synergy effectively.
@@ 2.2 Vision-language models and multi-modal prompt learning
Recent research [64, 63, 2, 3, 48, 47, 49] highlights prompt learning as a viable alternative to VLM fine-tuning on downstream tasks. To this end, unlike uni-modal approaches for CLIP, multi-modal deep prompts synergize its visual and textual components. [20] proposed to learn both visual and textual prompts and showed that initializing the visual prompts from the textual counterparts enhances the performance. In addition, [21] and [37] introduced regularization and feature consistency to prevent overfitting and ensure textual variety. However, existing methods mainly employ unidirectional token sharing, restricting the overall generalizability and semantic depth of the learned embeddings. Contrarily, we propose a bilateral approach to disseminate relevant cross-modal insights across CLIP’s branches, establishing a more effective multi-modal prompting paradigm.
@@ 3 Methodology
In the following, we delve into the initialization of text inputs and provide a detailed explanation of the image-text embeddings of CLIP in Section 3.1. Moving forward, our image-driven textual prompting approach is elaborated upon in Section 3.2, while Section 3.3 addresses the proposed visual prompting methodology. For fine-grained sketch-photo feature association, we discuss the proposed conditional cross-modality jigsaw task and the related details in Section 3.4. Finally, in Section 3.5, we mention the considered loss objectives. A list of important variables is summarized in the Supplementary.
@@ 3.1 Initialization of visual-textual embeddings
The pre-trained CLIP model operates on two modalities: text and image. It consists of a transformer-based [53] text encoder ($\mathcal{F}_{t}$) and ViT-based [9] image encoder ($\mathcal{F}_{v}$). They both contain $\mathcal{L}$ transformer encoder layers. $\mathcal{F}_{t}$ generates feature representations for text descriptions to capture semantic information. Initially, it tokenizes the text inputs, consisting of $\mathcal{J}$ words, and projects them into word embeddings $\mathbf{W}_{0}=[w^{1}_{0},w^{2}_{0},\cdots,w^{\mathcal{M}}_{0}]\in\mathbb{R}^{\mathcal{M}\times d_{t}}$, where [.,.][.,.][ . , . ] denotes stacking and concatenation, $\mathcal{M}$ refers the number of embedding tokens ($77$ per class-prompt) and $d_{t}$ is the dimension of text tokens. In our approach, we use the input texts as ‘‘sketch/photo of a [CLS]’’ for sketches and photos, respectively, for (G)ZS-SBIR, meaning $\mathcal{J}=5$. Here, the [CLS] token represents class embeddings, completing the prompt embedding $\mathbf{W}_{0}$. For the FG-ZS-SBIR task, we use common text input, ‘‘visual representation of [CLS]’’, for both sketches and photos. We obtain the $l$-th layer embedding of $\mathcal{F}_{t}$, denoted as $\mathcal{F}_{t}^{l}$, as follows,
@@
$[\mathbf{c}_{l},\mathbf{E}_{l}]=\mathcal{F}_{v}^{l}([\mathbf{c}_{l-1},\mathbf{E}_{l-1}])\in\mathbb{R}^{1+\mathcal{N}\times d_{v}}\qquad l=1,2,\cdots,\mathcal{L}\vspace{-0.1cm}$ (2)
@@ 3.2 Proposed vision-guided deep textual prompting
[_$[\_,\mathbf{W}_{l}]=\mathcal{F}_{t}^{l}([\mathbf{T}_{l-1};\mathbf{W}_{l-1}])\in\mathbb{R}^{\mathcal{M}\times d_{t}}\qquad l=1,2,\cdots,\mathcal{L}\vspace{-0.1cm}$ (3)
@@ 3.3 Proposed text-guided deep visual prompting
In addition, we enforce token sharing from each layer of $\mathcal{F}_{t}$ to the respective layer input of $\mathcal{F}_{v}$. Note that, from $l\geq 1$, $\mathbf{W}_{l}$ implicitly includes visual knowledge as per our proposed textual prompting. Secondly, as opposed to the visual prompting approach proposed in [20], which only shares the learnable textual tokens with the visual branch, we propose to consider all the tokens from $\mathbf{W}_{l}$ over all the classes present in $\mathcal{C}^{s}$ to be included into the input space of $\mathcal{F}_{v}^{l}$. This class-agnostic knowledge-sharing approach helps diminish the semantic gap in the learned embeddings.
@@
$[\mathbf{c}_{l},\_;\mathbf{E}_{l}]=\mathcal{F}_{v}^{l}([\mathbf{c}_{l-1},\mathbf{V}_{l-1};\mathbf{E}_{l-1}])\in\mathbb{R}^{1+\mathcal{N}\times d_{v}}\qquad l=1,2,\cdots,\mathcal{L}\vspace{-0.1cm}$ (5)
@@ 3.4 Proposed conditional cross-modal jigsaw solver for fine-grained sketch-photo feature association
To address these challenges, our proposed method incorporates positive and negative counterparts ($p_{a}^{+},p_{a}^{-}$) from set $\mathcal{P}^{s}$ when resolving the jigsaw puzzle for a sample $s_{a}\in\mathcal{S}^{s}$ given its permuted version $s_{a}^{\prime}$ which is obtained through a permutation function $\delta$, given the random permutation $y^{\text{perm}}\in\mathcal{Y}^{\text{perm}}$: $s^{\prime}_{a}=\delta(s_{a},y^{\text{perm}})$. Here, the information embedded in $(s^{\prime}_{a},p_{a}^{+})$ and $(s^{\prime}_{a},p_{a}^{-})$ is then processed through a transformer-based jigsaw-solver network, $\mathcal{F}_{js}$, operating atop $\mathcal{F}_{v}$, to leverage $p_{a}^{+}$ to inform better the prediction of the permutation arrangements of $s^{\prime}_{a}$ compared to $p_{a}^{-}$, achieved through a hinge objective. Collectively, $\mathcal{F}_{js}$ directly resolves the jigsaw puzzle for $s^{\prime}_{a}$ when coupled with $s_{a}$. These combined losses are called $\mathcal{L}_{cjs}$. For simplicity, we define the fused inputs from pairs of images to $\mathcal{F}_{js}$ as follows: $r=[\mathcal{F}_{v}(s_{a}),\mathcal{F}_{v}(s^{\prime}_{a})]$, $r^{+}=[\mathcal{F}_{v}(p_{a}^{+}),\mathcal{F}_{v}(s^{\prime}_{a})]$, and $r^{-}=[\mathcal{F}_{v}(p_{a}^{-}),\mathcal{F}_{v}(s^{\prime}_{a})]$, respectively.